OpenStudy (anonymous):

Evaluate cos220cos40 + sin220sin40

OpenStudy (anonymous):

are you looking for the answer in terms of radians, degrees, or just reagular numbers?

OpenStudy (anonymous):

I honestly don't know. It doesn't say. Probably either degrees or regular numbers.

OpenStudy (anonymous):

220 - 40 = 180, that's gotta be a part of it, doncha think?

OpenStudy (anonymous):

thats probable

OpenStudy (anonymous):

ok man look in the back of ur book for questions next to this

OpenStudy (anonymous):

see what its asking for

OpenStudy (anonymous):

okay, cos(-40) = cos(40) and cos(a+b) = cos(a)cos(b) - sin(a)sin(b)...

OpenStudy (anonymous):

... so (a+b) = 220 + -40

OpenStudy (anonymous):

ohhh. thank you so much! I missed class today and my teacher just gave me the notes so I was a little confused! Thank you!

OpenStudy (anonymous):

not to be petty, but my friend and I are competing for medals...